# Joint identification of spatially variable genes via a network-assisted Bayesian regularization approach

### v1

@@ 5 Discussion
In this article, we have presented a novel Bayesian regularization approach for the joint identification of SV genes, with the network structure among genes accommodated. The proposed approach has two main advantages. First, attributing to the Bayesian framework, it can automatically incorporate the network structure through the Gaussian Graph Laplacian priors. Compared to most of the existing methods, the utilization of such network information provides more opportunities to search for more biologically sensible SV genes. Second, to tackle the confounding effects of cell type mixtures within spots, cell-type-specific parameters have been introduced to model the variations induced by diverse latent cellular compositions, which are not well accommodated by most existing methods. The extensive simulation studies have revealed that the proposed approach can achieve better performance. The biological implications of the findings from the application to the TNBC dataset have further supported the utility of the proposed approach.
We have considered a series of presupposed spatial patterns and introduced the parametric model. It can be of interest to extend the proposed framework to accommodate unknown spatial patterns using nonparametric functional techniques. The proposed approach has focused on the identification of SV genes. It has the potential to be extended to additionally accommodate region segmentation of the tissue section. In data analysis, we have utilized the PPI information for network construction. Other resources, such as Kyoto Encyclopedia of Genes and Genomes (KEGG) and High-quality INTeractomes (HINT) databases, can also be adopted. Furthermore, a literature search and GO enrichment analysis have demonstrated the significant implications of the findings. A more definitive confirmation from functional validations may be needed.
@@
Supplementary Materials for Joint identification of spatially variable genes via a network-assisted Bayesian regularization approach
@@
$\boldsymbol{\gamma}^{*}=$ $\sqrt{1-\tau_{\gamma}^{2}}\boldsymbol{\gamma}+\left(1-\sqrt{1-\tau_{\gamma}^{2}}\right)\sigma_{\gamma}^{2}\left(\mathbf{L}+\varepsilon\mathbf{I}_{2p}\right)^{-1}\nabla_{\boldsymbol{\gamma}}\log f(\mathbf{Y}\mid\boldsymbol{\gamma},\boldsymbol{-})$
$+\tau_{\gamma}N\left\{\boldsymbol{0},\sigma_{\gamma}^{2}\left(\mathbf{L}+\varepsilon\mathbf{I}_{2p}\right)^{-1}\right\}$
@@
$\nabla_{\gamma^{(d)}_{j}}\log f(\mathbf{Y}\mid\boldsymbol{\gamma},\boldsymbol{-})=\sum_{\left\{i:r_{ij}=0\right\}}\frac{\phi\left(Y_{ij}-c_{i}\mu_{ij}\right)}{\mu_{ij}\left(c_{i}\mu_{ij}+\phi\right)}\cdot\mu_{ij}\cdot\mathcal{K}\left(x_{id}\right)\cdot\nabla_{\gamma^{(d)}_{j}}\left\{\gamma^{(d)}_{j}\operatorname{I}\left(\left|\gamma^{(d)}_{j}\right|>\lambda\cdot\rho^{(d)}_{j}\right)\right\}.$
For the derivative of the hard-thresholding function, a smooth approximation is introduced that $\operatorname{I}\left(\left|\gamma_{j}^{(d)}\right|>\lambda\cdot\rho^{(d)}_{j}\right)\simeq\frac{1}{2}\left\{1+\frac{2}{\pi}\arctan\left(\frac{\left(\gamma_{j}^{(d)}\right)^{2}-\left(\lambda\cdot\rho^{(d)}_{j}\right)^{2}}{\epsilon}\right)\right\}$ for $\epsilon\rightarrow 0$, leading to
$\nabla_{\gamma^{(d)}_{j}}\left\{\gamma^{(d)}_{j}\operatorname{I}\left(\left|\gamma^{(d)}_{j}\right|>\lambda\cdot\rho^{(d)}_{j}\right)\right\}\simeq\operatorname{I}\left(\left|\gamma^{(d)}_{j}\right|>\lambda\cdot\rho^{(d)}_{j}\right)+\gamma^{(d)}_{j}\cdot\frac{2\gamma^{(d)}_{j}/\epsilon}{\pi\left\{1+\left(\left(\gamma^{(d)}_{j}\right)^{2}-\left(\lambda\cdot\rho^{(d)}_{j}\right)^{2}\right)^{2}/\epsilon^{2}\right\}},$
@@
Regarding the hyperparameter set, we assume $a_{\pi}=b_{\pi}=1$ for a weakly informative prior on $\pi_{j}$. We set $a_{\phi}=10$ and $b_{\phi}=0.1$ for $\text{Ga}(a_{\phi},b_{\phi})$, and $a_{\gamma}=3.5$ and $b_{\gamma}=0.5$ for $\operatorname{IG}\left(a_{\gamma},b_{\gamma}\right)$. $\sigma_{0j}^{2}$’s and $\sigma_{\alpha_{k}}^{2}$’s are set as $3^{2}$ for all $j$ and $k$ to yield vague priors for the baseline expression levels and cell-type-specific effects. We set $\lambda_{l}=0$ and $\lambda_{u}$ as the 90% quantile of $\left|\tilde{\boldsymbol{\beta}}\right|$ to help against false positive detection as suggested in Song et al. (2021). The sampling variances $\tau_{\mu_{0}}^{2},\tau_{\alpha}^{2},\tau_{\phi}^{2},\tau_{\gamma}^{2},$ and $\tau_{\lambda}^{2}$ are all adaptively chosen by tuning acceptance rates to 30% in simulation studies and 15% in real data analysis. We perform 2,000 iterations with the first 1000 discarded as burn-in and thinning by 10 in the simulations, and 3,000 iterations with the first 2000 burn-in and thinning by 10 in the data analysis.
S3 Additional simulation settings and results
S3.1 Detailed Settings for Basic Simulations
Figure S3.1: Partitioned spot regions considered in the basic simulations introduced in Section 3 of the main text. Regions are denoted by different colors, where the cellular compositions $\boldsymbol{w}_{i}$’s are independently sampled from Dirichlet distributions $\operatorname{Dirc}(1,1,1,1,1,1)$ (Region 1), $\operatorname{Dirc}(3,5,7,9,11,13)$ (Region 2), and $\operatorname{Dirc}(18,16,14,12,10,8)$ (Region 3).
@@
Detailed settings for the spatial effects:
  1. (a)
    Star network
@@
    For each informative sub-network, the transcription factor (TF) gene and its connected target genes have the same sign of spatial effect (either all positive or all negative), while the magnitudes satisfy that $\frac{\left|\beta_{\operatorname{TF}}^{(d)}\right|}{\sqrt{n_{\operatorname{target}}}}>\left|\beta_{\operatorname{target}}^{(d)}\right|$ with $n_{\operatorname{target}}$ being the number of the target genes regulated by TF. Specifically, the values of $\beta^{(1)}_{\operatorname{TF}}$ and $\beta^{(2)}_{\operatorname{TF}}$ are fixed as $1$ and $-1$ for all three patterns, while $\beta^{(1)}_{\operatorname{target}}=0.15$ and $\beta^{(2)}_{\operatorname{target}}=-0.25$ for linear pattern, $\beta^{(1)}_{\operatorname{target}}=0.2$ and $\beta^{(2)}_{\operatorname{target}}=-0.3$ for periodic pattern, and $\beta^{(1)}_{\operatorname{target}}=0.4$ and $\beta^{(2)}_{\operatorname{target}}=-0.5$ for exponential pattern.
    For the non-informative sub-networks, $\beta^{(1)}_{\operatorname{TF}}=\beta^{(2)}_{\operatorname{TF}}=\beta^{(1)}_{\operatorname{target}}=\beta^{(2)}_{\operatorname{target}}=0$.
  2. (b)
    Scale-free network
@@
    For each informative sub-network, the spatial effects of the hub gene and its connected regular genes are randomly assigned as positive or negative. In addition, the magnitudes of $\beta^{(d)}_{\operatorname{hub}}(d=1,2)$ are simulated from $\operatorname{Unif}(1,1.2)$ for all three patterns. For the magnitudes of the regular SV genes, consider two types with large and small signal strengths, respectively. Specifically, for both $d=1,2$, the magnitudes are simulated as follows: $|\beta^{(d)}_{\operatorname{regular}}|\sim\operatorname{Unif}(a\cdot 0.1,a\cdot 0.2)$ for linear pattern, $|\beta^{(d)}_{\operatorname{regular}}|\sim\operatorname{Unif}(a\cdot 0.2,a\cdot 0.3)$ for periodic pattern, and $|\beta^{(d)}_{\operatorname{regular}}|\sim\operatorname{Unif}(a\cdot 0.4,a\cdot 0.5)$ for exponential pattern, where $a=1$ and 2 for small and large strengths, respectively.
    For the non-informative sub-networks, $\beta^{(1)}_{\operatorname{hub}}=\beta^{(2)}_{\operatorname{hub}}=\beta^{(1)}_{\operatorname{regular}}=\beta^{(2)}_{\operatorname{regular}}=0$.
S3.2 Implementation Details of the Competing Methods
  1. (a)
    SpatialDE: The source code for SpatialDE is publicly available on https://github.com/Teichlab/SpatialDE. Following the instructions, we filter the practically unobserved genes with total observations less than 3. The linear regression is conducted on the raw count matrix to account for the potential bias caused by library size or sequencing depth. The main function for SV gene detection is SpatialDE.run() with the arguments being the coordinate information $\left(x_{i1},x_{i2}\right)$ and the sample residual expressions corrected by the linear regression. To account for multiple testing, the genes with qval less than 0.05 are identified as SV.
  2. (b)
    SPARK: The source code for SPARK is publicly available on https://github.com/xzhoulab/SPARK. Following the instructions, we filter genes that are expressed in less than 10% spots and spots with the total observations less than 10. The statistical model under the null hypothesis is first fit by employing the function spark.vc(), and the function spark.test() is subsequently employed for SV gene detection. Then, the genes with adjusted_pvalue less than 0.05 are identified as SV.
  3. (c)
    SPARK-X: The source code for SPARK-X is publicly available on https://github.com/xzhoulab/SPARK. The SV gene detection is implemented with the main function sparkx() where the argument option is set as “mixture” for multiple kernels testing. To account for the impact of latent cellular compositions, as conducted in the original paper, each spot is assigned its major cell type, with the argument X_in being the matrix of binary indicators. Then, the genes with adjustedPval less than 0.05 are identified as SV.
  4. (d)
    MERINGUE: The source code for MERINGUE is publicly available on https://github.com/JEFworks-Lab/MERINGUE. The non-expressed genes and spots are filtered following the tutorial. Normalization is conducted for the raw count matrix by employing the function normalizeCounts(). For SV gene detection, the neighborhood relationships are first constructed using the function getSpatialNeighbors() with the argument filterDist set as the default value 2.5. Then, the SV genes are identified with the function filterSpatialPatterns() with the arguments minPercentCells set as 0.05 to restrict that the SV genes are driven by more than 5% of spots. The adjusted significance threshold is set as 0.05 through setting the arguments adjustPv as TRUE and alpha as 0.05.
  5. (e)
    CTSV and CTSV-g: The source code for CTSV is publicly available on https://github.com/jingeyu/CTSV. The SV gene detection is conducted through the main function ctsv(). Specifically, the cell-type-specific SV genes are identified with the argument $W$ being the $n\times K$ matrix composed of $w_{i}$’s ($i=1,\ldots,n$). The final SV gene set is the union of all cell-type-specific SV genes. We also conduct the global SV gene detection without accommodation for the cellular composition in the simulation studies (CTSV-g). Specifically, the argument $W$ is set as $(1,\ldots,1)^{\mathrm{T}}_{(n)}$. For both CTSV and CTSV-g, the SV genes are identified through function svGene() with the significance threshold thre.alpha set as 0.05.
  6. (f)
    HEARTSVG: The source code for HEARTSVG is publicly available on https://github.com/cz0316/HEARTSVG. The detection of SV genes is conducted through the main function heartsvg() with the raw count matrix first scaled as recommended. The Holm method is further conducted for multiple testing control. The genes with adjusted P values p_adj less than 0.05 are identified as SV.
S3.3 Additional Simulation Results under the Scenarios with a High Dropout Rate
@@
Table S3: Simulation results under the scenarios with a high dropout rate, where FDR (BFDR) is controlled to be $<$0.05. In each cell, mean (SD) is based on 50 replicates.
@@
Figure S3.3: Comparison boxplots of Recall for large (L:Recall) and small (S:Recall) signals, overall Recall, Precision, and F1 based on 50 replicates under scenarios with scale-free network and (A) linear pattern, (B) exponential pattern, and (C) periodic pattern, where FDR (BFDR) is controlled to be $<$0.05.
S3.4 Simulations under Model Misspecification
We considered two types of model misspecification scenarios as follows.
@@
    Poisson model considered in Sun et al. (2020). The raw count data is generated as $Y_{ij}\sim\operatorname{Poi}(N_{i}\cdot\lambda_{ij})$, where $\log\lambda_{ij}=\operatorname{Intercept}_{ij}+\epsilon_{ij}+\sum_{k=1}^{K}\alpha_{k}w_{ik}$ with $\operatorname{Intercept}_{ij}$ and $\epsilon_{ij}$ being the intercept and residual error term, respectively. Specifically, following Sun et al. (2020), consider 260 spots collected in the mouse olfactory bulb study (Ståhl et al., 2016) with the corresponding parameters inferred from the data analysis conducted by spatialDE and SPARK, and $p=5,000$ genes. The network is set as the same as (M1) (as shown in Figure S3.4(B)), except that all genes in the first ten sub-networks are SV, resulting in a total number of 500 SV genes. For non-SV genes, $\operatorname{Intercept}_{ij}$ is set to be $-10.2$ across all spots, which corresponds to the median of the intercept estimates in the mouse olfactory data analysis. For SV genes, the spots are categorized into two groups, including the group with low expression levels (green) and the group with high expression levels (pink), according to the spatial patterns of the identified SV genes in the mouse olfactory bulb data, as illustrated in Figure S3.5. In particular, $\operatorname{Intercept}_{ij}$’s for the low group are set as $-10.2$ while $\operatorname{Intercept}_{ij}$’s for the high group are set as $-8.8$ and $-17.8$ for the target and TF genes, respectively, according to the estimates for mean values. $\epsilon_{ij}$’s are independently simulated from $N(0,0.2^{2})$, which is approximately the first quantile of the non-spatial variance estimates. For cellular composition variations, consider $K=6$ underlying cell types, where $w_{ik}$’s of spots from high and low groups are independently sampled from $\operatorname{Dir}(3,5,7,9,11,13)$ and $\operatorname{Dir}(18,16,14,12,10,8)$, respectively. $\alpha_{k}$ is set as 0 for $k=1,\ldots,K$, so that the cellular composition variations do not contribute to the expression levels. $N_{i}$’s are obtained based on the real counts.
@@
Figure S3.4: (A) Partitioned spot regions considered in the model misspecification scenario (M1). Regions are denoted by different colors, where the cellular compositions $\boldsymbol{w}_{i}$’s are independently sampled from Dirichlet distributions $\operatorname{Dirc}(1,1,1,1,1,1)$ (Region 1), $\operatorname{Dirc}(1,3,5,7,9,11)$ (Region 2), $\operatorname{Dirc}(16,14,12,10,8,6)$ (Region 3), and $\operatorname{Dirc}(1,4,4,4,4,1)$ (Region 4). (B) Illustrative example of the simple regulatory network considered in model misspecification scenarios.
@@
Figure S3.6: Comparison boxplots of Recall, Precision, and F1 based on 50 replicates under (A) ZINB model considered in Yu and Luo (2022) and (B) Poisson model considered in Sun et al. (2020), where FDR (BFDR) is controlled to be $<$0.05.
@@
Figure S3.7: Comparison boxplots of Recall, Precision, and F1 based on 50 replicates, under (A) ZINB model considered in Yu and Luo (2022); (B) ZINB model considered in Section 3.1 of the main text with star network, linear spatial pattern, and low dropout rate = 0.1; (C) ZINB model considered in Section 3.1 of the main text with star network, linear spatial pattern, and high dropout rate = 0.5, where in each informative sub-network, about 20% of the genes connected to the TF genes are set as uninformative and FDR (BFDR) is controlled to be $<$0.05.

### v2

@@ 5 Discussion
In this article, we have presented a novel Bayesian regularization approach for the joint identification of SV genes, with the network structure among genes accommodated. The proposed approach has two main advantages. First, attributing to the Bayesian framework, it can automatically incorporate the network structure through the Gaussian Graph Laplacian priors. Compared to most of the existing methods, the utilization of such network information provides more opportunities to search for more biologically sensible SV genes. Second, to tackle the confounding effects of cell type mixtures within spots, cell-type-specific parameters have been introduced to model the variations induced by diverse latent cellular compositions, which are not well accommodated by most existing methods. The extensive simulation studies have revealed that the proposed approach can achieve better performance. The biological implications of the findings from the application to the Visium TNBC, Visium PLC, and Xenium PAC datasets have further supported the utility of the proposed approach.
The proposed approach has used a zero-inflated model to handle the excess zeros often seen in ST data, proving its necessity and effectiveness in our simulations. However, we recognize prior research (Zhao et al., 2022) indicating that modeling zero inflation is not always essential. Inspired by this, a valuable future direction is to systematically explore alternative zero-inflation handling strategies, balancing modeling accuracy and computational efficiency. We have identified SV genes by checking if their corresponding spatial coordinate-specific regression coefficients are nonzero. Though our approach uses regularization to induce sparsity, we acknowledge the value of inference-based alternatives relying on statistical testing, which we will explore in the future. We have considered preset spatial patterns and introduced a parametric model. It would be interesting to expand the proposed framework to accommodate unknown spatial patterns using nonparametric methods. Moreover, Su et al. (2025) noted that mean-value-based strategies may suffer from a lack of rotation invariance, wherein rotations alter spatial coordinates and consequently influence effect estimation and SV gene identification. We have carried out an exploratory analysis in Supplementary Section S8 (Wu et al., 2025) to evaluate the robustness of our method against spatial coordinate rotations. The results indicate that incorporating prior network information can partially mitigate the sensitivity of our approach to rotation. However, inconsistent SV gene detection persists across different rotation scales. Further investigation integrating rotation-invariant strategies appears promising, such as constructing predictors based on relative distances as suggested in Su and Cui (2025).
In our real-data analysis, we have employed Redeconve (Zhou et al., 2023) to estimate cell-type proportions before detecting SV genes. This is a widely-used strategy in existing studies (Zhu, Sun and Zhou, 2021; Yu and Luo, 2022; Cable et al., 2022b) and has demonstrated satisfactory performance with biological significance. Although Redeconve relies on expression data, the additional network information we have incorporated can mitigate this dependency to some extent. Furthermore, our simulation studies (Section 3.6) indicate that the proposed method is quite robust to noise in cell-type proportion estimates. Exploring alternative cell-type estimation approaches is a potential avenue for future work. In addition to cellular composition information, we utilized PPI information for network construction. Other resources, such as the Kyoto Encyclopedia of Genes and Genomes (KEGG) and High-quality INTeractomes (HINT) databases, could also be used. Additionally, as an alternative, the network could be directly estimated from the analyzed data. A literature search and GO enrichment analysis have demonstrated the significant implications of the findings. A more definitive confirmation from functional validations may be needed.
{acks}
[Acknowledgments] The authors would like to thank the anonymous referees, an Associate Editor, and the Editor for their constructive comments which led to a significant improvement of this article. Mingcong Wu and Yang Li are joint first authors, and Mengyun Wu is the corresponding author. {funding} This research was supported by the National Natural Science Foundation of China (12071273); Shanghai Rising-Star Program (22QA1403500); Shanghai Science and Technology Development Funds (23JC1402100); Shanghai Research Center for Data Science and Decision Technology; Public Health &\& Disease Control and Prevention, the MOE Project of Key Research Institute of Humanities and Social Sciences (22JJD910001); National Institutes of Health (CA204120 and CA121974); and National Science Foundation (2209685).
[…]
@@
Supplement to “Joint identification of spatially variable genes via a network-assisted Bayesian regularization approach”
@@
$\boldsymbol{\gamma}^{*}=\sqrt{1-\tau_{\gamma}^{2}}\boldsymbol{\gamma}+\left(1-\sqrt{1-\tau_{\gamma}^{2}}\right)\sigma_{\gamma}^{2}\left(\mathbf{L}+\varepsilon\mathbf{I}_{2p}\right)^{-1}\nabla_{\boldsymbol{\gamma}}\log f(\mathbf{Y}\mid\boldsymbol{\gamma},\boldsymbol{-})+\tau_{\gamma}N\left\{\boldsymbol{0},\sigma_{\gamma}^{2}\left(\mathbf{L}+\varepsilon\mathbf{I}_{2p}\right)^{-1}\right\}$
@@
$\nabla_{\gamma^{(d)}_{j}}\log f(\mathbf{Y}\mid\boldsymbol{\gamma},\boldsymbol{-})=\sum_{\left\{i:r_{ij}=0\right\}}\frac{\phi\left(Y_{ij}-c_{i}\mu_{ij}\right)}{\mu_{ij}\left(c_{i}\mu_{ij}+\phi\right)}\,\mu_{ij}\,\mathcal{K}\left(x_{id}\right)\,\nabla_{\gamma^{(d)}_{j}}\left\{\gamma^{(d)}_{j}\operatorname{I}\left(\left|\gamma^{(d)}_{j}\right|>\lambda\,\rho^{(d)}_{j}\right)\right\}$.
For the derivative of the hard-thresholding function, a smooth approximation is introduced that $\operatorname{I}\left(\left|\gamma_{j}^{(d)}\right|>\lambda\rho^{(d)}_{j}\right)\simeq\frac{1}{2}\left\{1+\frac{2}{\pi}\arctan\left(\frac{\left(\gamma_{j}^{(d)}\right)^{2}-\left(\lambda\rho^{(d)}_{j}\right)^{2}}{\epsilon}\right)\right\}$ for $\epsilon\rightarrow 0$, leading to
$\nabla_{\gamma^{(d)}_{j}}\left\{\gamma^{(d)}_{j}\operatorname{I}\left(\left|\gamma^{(d)}_{j}\right|>\lambda\rho^{(d)}_{j}\right)\right\}\simeq\operatorname{I}\left(\left|\gamma^{(d)}_{j}\right|>\lambda\rho^{(d)}_{j}\right)+\gamma^{(d)}_{j}\frac{2\gamma^{(d)}_{j}/\epsilon}{\pi\left\{1+\left(\left(\gamma^{(d)}_{j}\right)^{2}-\left(\lambda\rho^{(d)}_{j}\right)^{2}\right)^{2}/\epsilon^{2}\right\}},$
@@
[…]
S3 Computation efficiency
We assess the computational efficiency of our proposed method using both simulated and real datasets. For simulated datasets, with a fixed gene size of 5,000 and 2000 MCMC iterations, the average computer times for spot sizes of 1,000, 5,000, 10,000, and 50,000 are 0.37, 1.53, 2.82, and 15.84 hours, respectively. In real data analysis, the computer times for a full 3,000 iteration MCMC run (which is sufficient to ensure convergence) are 1.2, 2.7, and 5.3 hours for the Visium TNBC (1,108 spots with 5,000 genes), Visium PLC (3,181 spots with 5,000 genes), and Xenium PAC (164,274 single cells with 474 genes) datasets, respectively. All analyses are implemented on a MacBook Pro computer with Intel Core i5, 4 cores, and 16 GB RAM, using 8 threads for paralleling. These results demonstrate that our proposed approach is computationally feasible even for large-scale cell-resolution ST data (mostly with a few genes).
S4 Detailed simulation settings and additional simulation results
S4.1 Graphical representation for cellular compositions and network structures in the basic simulations
Figure S4.1: Partitioned spot regions considered in the basic simulations introduced in Section 3.1 of the main text. Regions are denoted by different colors, where the cellular compositions $\boldsymbol{w}_{i}$’s are independently sampled from Dirichlet distributions Dirc(1,1,1,1,1,1)(Region 1), Dirc(3,5,7,9,11,13) (Region 2), and Dirc(18,16,14,12,10,8) (Region 3).
@@
S4.2 Detailed settings for spatial effects in the basic simulations
The spatial effects associated with the two specified network structures are set as follows.
  1. 1.
    Star network:
@@
    Denote $\beta_{\operatorname{TF}}$ and $\beta_{\operatorname{target}}$ as the spatial effect parameters for the transcription factor (TF) gene and connected target gene, respectively. For the non-informative sub-networks, $\beta^{(1)}_{\operatorname{TF}}=\beta^{(2)}_{\operatorname{TF}}=\beta^{(1)}_{\operatorname{target}}=\beta^{(2)}_{\operatorname{target}}=0$. For each informative sub-network, the transcription factor (TF) gene and its connected target genes have the same sign of spatial effect (either all positive or all negative), while the magnitudes satisfy that $\frac{\left|\beta_{\operatorname{TF}}^{(d)}\right|}{\sqrt{n_{\operatorname{target}}}}>\left|\beta_{\operatorname{target}}^{(d)}\right|$ with $n_{\operatorname{target}}$ being the number of the target genes regulated by TF gene. The specified values of $\beta_{\operatorname{TF}}$ and $\beta_{\operatorname{target}}$ for informative sub-networks under different spatial patterns are detailed in Table S3.
  2. 2.
    Scale-free network:
@@
    Denote $\beta_{\operatorname{Hub}}$ and $\beta_{\operatorname{Regular}}$ as the spatial effect parameters for the hub gene and connected regular gene, respectively. For the non-informative sub-networks, $\beta^{(1)}_{\operatorname{Hub}}=\beta^{(2)}_{\operatorname{Hub}}=\beta^{(1)}_{\operatorname{Regular}}=\beta^{(2)}_{\operatorname{Regular}}=0$. For each informative sub-network, the spatial effects of the hub gene and its connected regular genes are randomly assigned as positive or negative. In addition, the magnitudes of $\beta^{(d)}_{\operatorname{Hub}}(d=1,2)$ are simulated from $\operatorname{Unif}(\operatorname{Hub}_{a},\operatorname{Hub}_{b})$, while the magnitudes of $\beta^{(d)}_{\operatorname{Regular}}(d=1,2)$ are randomly generated from $\operatorname{Unif}(\operatorname{Regular}_{a},\operatorname{Regular}_{b})$ or $\operatorname{Unif}(2\operatorname{Regular}_{a},2\operatorname{Regular}_{b})$ with the probability generated from a $\operatorname{Bern}(0.5)$ distribution, representing large and small signals, respectively. The specified values of these parameters for informative sub-networks under different spatial patterns are detailed in Table S3.
Table S3: Detailed values of spatial effects for informative sub-networks in the basic simulations.
Network Spatial pattern Signal strength
Star Linear $\beta^{(1)}_{\operatorname{TF}}=1,\beta^{(2)}_{\operatorname{TF}}=-1$ $\beta^{(1)}_{\operatorname{target}}=0.15,\beta^{(2)}_{\operatorname{target}}=-0.25$
Exponential $\beta^{(1)}_{\operatorname{target}}=0.4,\beta^{(2)}_{\operatorname{target}}=-0.5$
Periodic $\beta^{(1)}_{\operatorname{target}}=0.2,\beta^{(2)}_{\operatorname{target}}=-0.3$
Scale-free Linear $\text{Hub}_{a}=1.0,\text{Hub}_{b}=1.2$ $\text{Regular}_{a}=0.1,\text{Regular}_{b}=0.2$
Exponential $\text{Regular}_{a}=0.3,\text{Regular}_{b}=0.4$
Periodic $\text{Regular}_{a}=0.2,\text{Regular}_{b}=0.3$
S4.3 Additional results for the basic simulations
@@
Table S4: Simulation results under the scenarios with a high dropout rate, where FDR (BFDR) is controlled to be <<0.05. In each cell, mean (SD) is based on 50 replicates.
@@
S4.4 Results for the SV-free simulated datasets
Figure S4.3: Comparison boxplots of False Positive Rate (FPR) based on 50 replicates under SV-free simulation scenarios, where FDR(BFDR) is controlled to be <<0.05.
S4.5 Results for the validation of zero-inflation distribution
Figure S4.4: Comparison boxplots of Recall, Precision, and F1 based on 50 replicates under scenarios with star network and periodic pattern, where FDR (BFDR) is controlled to be <<0.05.
S4.6 Results for the validation of network assistance strategy
Figure S4.5: Comparison boxplots of Recall for large (L:Recall) and small (S:Recall) signals, overall Recall, Precision, and F1 based on 50 replicates under scenarios with scale-free network and (A) linear pattern, (B) exponential pattern, and (C) periodic pattern, where FDR (BFDR) is controlled to be <<0.05.
S4.7 Detailed settings for model misspecified simulations and results
We consider three types of model misspecification scenarios as follows.
@@
    The generalized linear spatial model where the spatial variability is introduced through the Gaussian covariance matrix as adopted in Svensson, Teichmann and Stegle (2018) and Sun, Zhu and Zhou (2020). Specifically, for gene $j$, the raw count data is generated through a zero-inflated Poisson distribution $y_{j}\left(x_{i}\right)\sim\operatorname{Poi}\left(\lambda_{j}\left(x_{i}\right)\right)^{(1-r_{ij})}\delta_{0}(y_{j}\left(x_{i}\right))^{r_{ij}}$ with $\left(\log\left(\lambda_{j}\left(x_{1}\right)\right),\ldots,\log\left(\lambda_{j}\left(x_{n}\right)\right)\right)^{\mathrm{T}}\sim\mathrm{N}\left(\mu_{j}\mathbf{1},\tau_{j}^{2}\Sigma+\sigma_{j}^{2}\mathbf{I}\right)$. Consider $n=400$ spots located on a $20\times 20$ grid and $p=5,000$ genes. The network is set as the same as (M1) (as shown in Figure S4.6(B)). The genes in the first ten sub-networks are SV, resulting in a number of 500 SV genes. Following Sun, Zhu and Zhou (2020), the Gaussian covariance matrix $\Sigma$ is set as $(\Sigma_{ii^{\prime}})_{n\times n}=\left(\exp\left(-\frac{\left\|x_{i}-x_{i^{\prime}}\right\|^{2}}{2l^{2}}\right)\right)_{n\times n}$ with length scale $l$ set as 1. For SV gene, $\tau^{2}_{j}$ (accounts for the expression variance attributable to spatial effects) is set as 0.7 while $\sigma_{j}^{2}$ is set as 0.3. For non-SV gene, $\tau^{2}_{j}=0$ and $\sigma^{2}_{j}=1$. No cellular composition variations are considered in this setting. ${r}_{ij}$’s are simulated from a Bernoulli distribution $\operatorname{Bern}\left(0.5\right)$.
  3. (M3)
@@
    Consider 260 spots collected in the mouse olfactory bulb study (Ståhl et al., 2016) and generate simulated data based on Poisson generalized linear model adopted in Sun, Zhu and Zhou (2020) with the corresponding parameters inferred from the data analysis results conducted by spatialDE and SPARK. Specifically, following the setting of Sun, Zhu and Zhou (2020), the raw count data is generated as $Y_{ij}\sim\operatorname{Poi}(N_{i}\cdot\lambda_{ij})$, where $\log\lambda_{ij}=\operatorname{Intercept}_{ij}+\epsilon_{ij}+\sum_{k=1}^{K}\alpha_{k}w_{ik}$ with $\operatorname{Intercept}_{ij}$ and $\epsilon_{ij}$ being the intercept and residual error term, respectively. Consider $p=5,000$ genes, the network is set as the same as (M1) and (M2) (Figure S4.6(B)) with all genes in the first ten sub-networks set as SV. For non-SV genes, $\operatorname{Intercept}_{ij}$ is set to be 10.2-10.2 across all spots, which corresponds to the median of the intercept estimates in the mouse olfactory data analysis. For SV genes, the spots are categorized into two groups, including the group with low expression levels (green) and the group with high expression levels (pink), according to the spatial patterns of the identified SV genes in the mouse olfactory bulb data, as illustrated in Figure S4.7. In particular, $\operatorname{Intercept}_{ij}$’s for the low group are set as 10.2-10.2 while $\operatorname{Intercept}_{ij}$’s for the high group are set as 8.8-8.8 and 17.8-17.8 for the target and TF genes, respectively, according to the estimates for mean values. $\epsilon_{ij}$’s are independently simulated from $\mathrm{N}(0,0.2^{2})$, which is approximately the first quantile of the non-spatial variance estimates. For cellular composition variations, consider $K=6$ underlying cell types, where $w_{ik}$’s of spots from high and low groups are independently sampled from $\operatorname{Dirc}(3,5,7,9,11,13)$ and $\operatorname{Dirc}(18,16,14,12,10,8)$, respectively. $\alpha_{k}$ is set as 0 for $k=1,\ldots,K$, so that the cellular composition variations do not contribute to the expression levels. $N_{i}$’s are obtained based on the real counts.
@@
Figure S4.6: (A) Partitioned spot regions considered in the model misspecification scenario (M1). Regions are denoted by different colors, where the cellular compositions $\boldsymbol{w}_{i}$’s are independently sampled from Dirichlet distributions Dirc(1,1,1,1,1,1) (Region 1), Dirc(1,3,5,7,9,11) (Region 2), Dirc(16,14,12,10,8,6) (Region 3), and Dirc(1,4,4,4,4,1) (Region 4). (B) Illustrative example of the simple regulatory network considered in model misspecification scenarios.
@@
Figure S4.8: Comparison boxplots of Recall, Precision, and F1 based on 50 replicates in model misspecification scenarios (A) M1, (B) M2, and (C) M3, where FDR (BFDR) is controlled to be <<0.05.
S4.8 Results for the examination on the noise of network and cellular composition
@@
Figure S4.9: Comparison boxplots of Recall, Precision, and F1 based on 50 replicates, under (A) ZINB model considered in Yu and Luo (2022); (B) ZINB model considered in Section 3.1 of the main text with star network, linear spatial pattern, and low dropout rate = 0.1; (C) ZINB model considered in Section 3.1 of the main text with star network, linear spatial pattern, and high dropout rate = 0.5, where in each informative sub-network, about 20% of the genes connected to the TF genes are set as uninformative and FDR (BFDR) is controlled to be <<0.05.